5. Find the general solution to y – 4y + 4y = - +1+sin z. 6. Find the general solution to y – 2y + y = tel 1+2 7. Find the

please solve problems 5, 6, and 7. The last guy didn't answer all three and I need all three. If you're gonna answer at least do 6 and 7 since we already have 5.
Answer #1

و :a م ن تم ;م Ain ر% = با + اوا - + و + روا دا ده ابن عطر و دو + m 4m_ | ود 2ر2 - m) 2 عره +ع) = م ( +( 2 ) ا +70-236 + ozP1 = 4(42+ & J + fakinn + 4casa 3. - General son of Cifferential equ į Scannedin (G+€43.0* + [détrú$]+=> [3drism + 4 com] rs